OpenStudy (anonymous):

How can I solve this equation without making the algebra tedious. Plus I don't want to use the reduction formula

OpenStudy (anonymous):

\[y'' + 9 y = 25 x \cos(2x)\]

OpenStudy (anonymous):

\[y_{p} = (Ax+b)\cos(2x) + (Cx+d)\sin(2x)\]

OpenStudy (anonymous):

Is there a technique to make the simplification easy, it turns out that whenever I try to do this equation, I end up with lines of variables and I am not sure If I can go wrong anywhere

OpenStudy (anonymous):

@pooja195 @Michele_Laino